OpenStudy (anonymous):

Explain how you can just look at an equation or an expression and correctly identify which type of change it represents, growth or decay.

hero (hero):

Let r be any rate in decimal form: \(y = a(1+r)^t\) will represent growth \(y = a(1 - r)^t\) will represent decay

hero (hero):

For example, if the rate was 12%,then r = 0.12 and \(y = a(1.12)^t\) would represent a growth rate of 12% \(y = a(0.88)^t\) would represent a decay factor of 12%

OpenStudy (anonymous):

May you please explain it?

hero (hero):

For more info, read this: http://www.regentsprep.org/regents/math/algebra/ae7/expdecayl.htm

OpenStudy (anonymous):

Thank you :) I get it now.